OpenStudy (anonymous):

Let f(x) = x2 – 7x + 3 and g(x) = x + 4. Find f(x) + g(x).

OpenStudy (anonymous):

set the two equal to each other and solve

OpenStudy (anonymous):

like so:

OpenStudy (anonymous):

x^2 – 7x + 3=x + 4

OpenStudy (anonymous):

dzhy e

OpenStudy (anonymous):

sory stupid cp froze!

OpenStudy (anonymous):

why would you set the two equal to each other...you are not solving for x...the question is just asking to add f(x) and g(x)

OpenStudy (anonymous):

the correct way to solve this problem is x^2-7x+3+x+4 =x^2-6x+7